OpenStudy (anonymous):

can someone help me: find x if y=-4 3y-(6/x)=12

OpenStudy (anonymous):

well it is 6 over x

OpenStudy (anonymous):

Replace y with -4 So, 3(-4)-(6/x)=12 -12-((6/x)=12 -(6/x)=24

OpenStudy (anonymous):

kk then what do you do

OpenStudy (anonymous):

because when i tried i got 6

OpenStudy (anonymous):